OpenStudy (sugarmochi):

Find the midpoint of the chord formed by the line x+y+4=0 and the hyperbola y=1/x

OpenStudy (3mar):

Could I help you?

OpenStudy (sugarmochi):

yeah

OpenStudy (3mar):

so What do you think? Any ideas?

OpenStudy (3mar):

Did you not try to plot these two function and notice the relationship between them graphically?

OpenStudy (3mar):

Like that form example? https://www.desmos.com/calculator/d8gcygp9qd

OpenStudy (sshayer):

x+y+4=0 y=1/x \[x \neq 0\] \[x+\frac{ 1 }{ x }+4=0,x^2+4x+1=0\] \[x=\frac{ -4\pm \sqrt{4^2-4*1*1} }{ 2*1 }=\frac{ -4 \pm \sqrt{12} }{ 2 }=-2 \pm \sqrt{3}\] therefore two points of intersection are \[x1=-2+\sqrt{3},x2=-2-\sqrt{3}\] mid point of x1 and x2 is \[x=\frac{( -2+\sqrt{3})+(-2-\sqrt{3}) }{ 2 }=-2\] x+y+4=0 \[when\ x=-2+\sqrt{3}\] \[-2+\sqrt{3}+y+4=0,y=-2-\sqrt{3}\] when \[x=-2-\sqrt{3},-2-\sqrt{3}+y+4=0,y=-2+\sqrt{3}\] point of intersections are \[\left( -2+\sqrt{3},-2-\sqrt{3} \right) \ and \ \left( -2-\sqrt{3},-2+\sqrt{3} \right)\] Let (x,y) be the mid point of these two points. then \[x=\frac{ x1+x2 }{ 2 },y=\frac{ y1+y2 }{ 2 }\] x co-ordinate of mid point i have calculated. similarly find y

OpenStudy (3mar):

Thank you for good explanation, @sshayer Really really nice....